OpenStudy (anonymous):

Simplify 3 logb m – 2 logb n. A. log b(m3n2) B. logb (m3/n2) C. logb (m2/n3) D. logb (m2n3)

OpenStudy (anonymous):

option B is correct

OpenStudy (anonymous):

\[3\log_b(m)-2\log_b(n)=\log_b(m^3)-\log_b(n^2)=\log_b(\frac{n^3}{n^2})\]

OpenStudy (anonymous):

I got B to, is that correct?

OpenStudy (anonymous):

ello?

OpenStudy (anonymous):

yes